what is 8 and 3/4 multipied by 3 and 7/8
Answer by stanbon(75887)

what is 8 and 3/4 multipied by 3 and 7/8?
---------
(35/4)*(31/8) = 1085/32 = 33 29/32
Cheers,
Stan H.